Opened 6 years ago

Last modified 4 years ago

#479 new flight processing

RG12/09, flight day 121/2013, Wessex

Reported by: knpa Owned by: besm
Priority: alpha 4 high Milestone:
Component: Processing: general Keywords:
Cc: Other processors:

Description (last modified by besm)

Data location: /users/rsg/arsf/arsf_data/2013/flight_data/unpacking/RG12_09-2013_121_Wessex

Data arrived from ARSF via network transfer on 13/05/2013

RG12/09 Scientific objective: Biodiversity and the provision of multiple ecosystem services in current and future lowland multifunctional landscapes

RG12/09 Priority: 8+

PI: France Gerard

Sensors:

Camera (Delivered 17/07/2013)
Eagle (Delivered 17/07/2013)
Hawk (Delivered 17/07/2013)
LiDAR (Delivered 17/07/2013)

Change History (27)

comment:1 Changed 6 years ago by besm

Starting nav. processing.

comment:2 Changed 6 years ago by besm

Finished nav. processing, started hyperspectral processing.

comment:3 Changed 6 years ago by besm

Starting LiDAR processing

comment:4 Changed 6 years ago by besm

SCTs for hyperspectral processing. The navigation was only valid for eagle lines 12, 13 and 15.

Eagle line 14 appeared to be positioned incorrectly, so doing its SCT would be impossible. It is not being processed, and eagle line 15 will be renamed to eagle line 14.

Flightline Eagle SCT Hawk SCT
1 2.05 1.73
2 1.20 0.87
3 2.82 2.42
4 2.04 -0.37
5 1.98 -0.41
6 2.68 0.28
7 3.03 0.67
8 3.00 0.68
9 2.41 0.02
10 3.79 1.41
11 2.63 0.23
12 -0.04 0.71
13 -0.05 0.70
14 -0.01 19.42

comment:5 Changed 6 years ago by knpa

  • Description modified (diff)

comment:6 Changed 6 years ago by besm

LiDAR pitch and roll errors

roll: -0.001132
pitch: -0.003185

These result in good results in many cases, but in the worst cases still results in ~40cm displacement between projections. I chose these pitch and roll errors because they tend to produce good results throughout the flight with the worse results evenly spread with projection displacements roughly balanced on either side of zero. Getting better results would require assigning unique pitch and roll values to every flightline, and the loss in accuracy is probably not worth it.

comment:7 Changed 6 years ago by besm

RCD delivery ready for delivery check. (Forgot to note when RCD processing was started. Last Friday I think.)

comment:8 Changed 6 years ago by knpa

  • Description modified (diff)

comment:9 Changed 6 years ago by knpa

  • Priority changed from alpha 4 medium to alpha 4 high

comment:10 Changed 6 years ago by anhi

RCD Delivery check:

No problems found.

comment:11 Changed 6 years ago by besm

Completed LiDAR delivery - ready for delivery check.

comment:12 Changed 6 years ago by besm

Completed hyperspectral delivery. Found that line 1 for both eagle and hawk was a (pretty useless) test line and seemed to interfere with making a good mosaic image, so I removed it based on Mark's advice.

The removed line can be found in the processing/line1/ folder, if needed.

comment:13 Changed 6 years ago by besm

Started copying LiDAR and RCD data to disk 162.

comment:14 Changed 6 years ago by anhi

Hyperspectral Delivery check:

  • Move the logsheet in.
  • Hawk line 13 sct looks like it could be improved.
  • Mention the hawk level one files have some underflow - mostly on the latter bands.

comment:15 Changed 6 years ago by edfi

Starting LiDAR D/C

comment:16 Changed 6 years ago by besm

It seems that I miscalculated the SCT for hawk line 13. Here's the updated SCT table.

Flightline Eagle SCT Hawk SCT
1 2.05 1.73
2 1.20 0.87
3 2.82 2.42
4 2.04 -0.37
5 1.98 -0.41
6 2.68 0.28
7 3.03 0.67
8 3.00 0.68
9 2.41 0.02
10 3.79 1.41
11 2.63 0.23
12 -0.04 0.71
13 -0.05 -0.68
14 ? 19.42
15 -0.01

comment:17 Changed 6 years ago by edfi

LiDAR D/C complete. All looks fine, ready to deliver. I agree with your previous comments about the Pitch and Roll values.

comment:18 Changed 6 years ago by besm

Created new delivery with updated SCT values. Ready for hyperspectral delivery check.

comment:19 Changed 6 years ago by anhi

Hyperspectral Delivery check:

  • Screenshots still have the SCT values in the names.

No other problems

comment:20 Changed 6 years ago by besm

Fixed. Please delivery check. (Do not run zip_mapped.sh, my machine is doing that already.)

comment:21 Changed 6 years ago by anhi

Hyperspectral Delivery recheck:

No problems found

comment:22 Changed 6 years ago by besm

Copying hyperspectral data to disk 161.

comment:23 Changed 6 years ago by besm

Sent hyperspectral data on disk 161, and camera and LiDAR on disk 162. Delivered to France Gerard at the Centre of Ecology Hydrology at OX108BB. 17 July 2013.

comment:24 Changed 6 years ago by besm

  • Description modified (diff)

comment:25 Changed 6 years ago by besm

Resent hyperspectral data on disk 168, and camera and LiDAR on disk 169 to France Gerard at same address. 08 August 2013.

comment:26 Changed 5 years ago by tipo

Split each hyperspectral and lidar line into 3 sections and redelivered.

comment:27 Changed 4 years ago by lah

Submitted to NEODC for archiving 16/09/2014.

Note: See TracTickets for help on using tickets.